Problem 1. Find the Maclaurin series for f(x) = 172 Find the interval of convergence for the Taylor series you found.
Answer #1

二o=1 0)=3 )=0 (4) = St-1) 9 10-23 = ギ =-340X2ョギ(0)=0 D4 fa=24CSx104DC0=242)! ⑥ 02495 キロ)=1+(H)xZ+CD XY toclearly it is only when a infinite 122 <1 I= [(-4,1)] Gol which converter & Tlak al ce